# ExpProof : Operationalizing Explanations for Confidential Models with ZKPs

### v2

@@ 1 Introduction
Explanations have been seen as a way to enhance trust in machine learning (ML) models by virtue of making them transparent. Although starting out as a debugging tool, they are now also widely proposed to prove fairness and sensibility of ML-based predictions for societal applications, in research studies (Langer et al., 2021; Smuha, 2019; Kästner et al., 2021; Von Eschenbach, 2021; Leben, 2023; Karimi et al., 2020; Wachter et al., 2017; Liao & Varshney, 2021) and regulations alike (Right to Explanation (Wikipedia contributors, 2025)). However, as discussed in detail by (Bordt et al., 2022), many of these use-cases are adversarial in nature where the involved parties have misaligned interests and are incentivized to manipulate explanations to meet their ends. For instance, a bank which denies loan to an applicant based on an ML model’s prediction has an incentive to return an incontestable explanation to the applicant rather than reveal the true workings of the model since the explanation can be used by the applicant to prove discrimination in the court of law (Bordt et al., 2022). In fact, many previous studies show that adversarial manipulations of explanations are possible in realistic settings with systematic and computationally feasible attacks (Slack et al., 2020; Shahin Shamsabadi et al., 2022; Slack et al., 2021; Yadav et al., 2024c). As such, despite the demand, explanations fail to be operational as a trust-enhancing tool.
@@ 4 Variants of LIME
BorderLIME. An important consideration for generating meaningful local explanations is that the sampled neighborhood should contain points from different classes (Laugel et al., 2018). Any reasonable neighborhood for an input far off from the decision boundary will only contain samples from the same class, resulting in vacuous explanations.
To remedy the problem, (Laugel et al., 2017, 2018) propose a radial search algorithm, which finds the closest point to the input $x$ belonging to a different class, $x_{border}$, and then uses $x_{border}$ as the input to LIME (instead of original input $x$). Their algorithm incrementally grows (or shrinks) a search area radially from the input point and relies on random sampling within each ‘ring’ (or sphere), looking for points with an opposite label. To cryptographically prove this algorithm, we would either have to reimplement the algorithm as-is or would have to give a probabilistic security guarantee (using a concentration inequality), both of which would require many classifier calls and thereby many proofs of inference, becoming inefficient in a ZKP system.
@@
Our algorithm samples $m$ directions and then starting from the original input $x$, takes $\delta$ steps until it finds a point with a different label along all these directions individually. The border point $x_{border}$ is that oppositely labeled point which is closest to the input $x$. Furthermore, unlike in the algorithm in (Laugel et al., 2017), our algorithm can exploit parallelization by searching along the different directions in parallel since these are independent.
@@ 8 Related Work
In the ML field ZKPs have been majorly used for verification of inferences made by models (Sun et al., 2024; Chen et al., 2024; Kang et al., 2022; Weng et al., 2023; Sun & Zhang, 2023; Feng et al., 2021; VI2, 2023; Lee et al., 2020; Zhang et al., 2020a; Liu et al., 2021; Singh et al., 2022; Fan et al., 2023). A line of work also focuses on proving the training of ML models using ZKPs (Burkhalter et al., 2021; Huang et al., 2022; Rückel et al., 2022; Garg et al., 2023; Abbaszadeh et al., 2024). More recently they’re also been used for verifying properties such as fairness (Yadav et al., 2024a; Shamsabadi et al., 2023; Toreini et al., 2023) and accuracy (Zhang et al., 2020b) of confidential ML models. Alternatively, (Waiwitlikhit et al., 2024) propose ZKPs as a general technique for auditing ML models broadly. Contrary to these and to the best of our knowledge, ours is the first work that identifies the need for proving explanations and provides ZKP based solutions for the same.

### v3

@@ 1 Introduction
Explanations have been seen as a way to enhance trust in machine learning (ML) models by virtue of making them transparent. Although starting out as a debugging tool, they are now also widely proposed to prove fairness and sensibility of ML-based predictions for societal applications, in research studies (Langer et al., 2021; Smuha, 2019; Kästner et al., 2021; Von Eschenbach, 2021; Leben, 2023; Karimi et al., 2022; Wachter et al., 2017; Liao & Varshney, 2021) and regulations alike (Right to Explanation (Wikipedia contributors, 2025)). However, as discussed in detail by (Bordt et al., 2022), many of these use-cases are adversarial in nature where the involved parties have misaligned interests and are incentivized to manipulate explanations to meet their ends. For instance, a bank which denies loan to an applicant based on an ML model’s prediction has an incentive to return an incontestable explanation to the applicant rather than reveal the true workings of the model since the explanation can be used by the applicant to prove discrimination in the court of law (Bordt et al., 2022). In fact, many previous studies show that adversarial manipulations of explanations are possible in realistic settings with systematic and computationally feasible attacks (Slack et al., 2020; Shahin Shamsabadi et al., 2022; Slack et al., 2021; Yadav et al., 2024c). As such, despite the demand, explanations fail to be operational as a trust-enhancing tool.
@@ 4 Variants of LIME
BorderLIME. An important consideration for generating meaningful local explanations is that the sampled neighborhood should contain points from different classes (Laugel et al., 2018b). Any reasonable neighborhood for an input far off from the decision boundary will only contain samples from the same class, resulting in vacuous explanations.
To remedy the problem, (Laugel et al., 2018a, b) propose a radial search algorithm, which finds the closest point to the input $x$ belonging to a different class, $x_{border}$, and then uses $x_{border}$ as the input to LIME (instead of original input $x$). Their algorithm incrementally grows (or shrinks) a search area radially from the input point and relies on random sampling within each ‘ring’ (or sphere), looking for points with an opposite label. To cryptographically prove this algorithm, we would either have to reimplement the algorithm as-is or would have to give a probabilistic security guarantee (using a concentration inequality), both of which would require many classifier calls and thereby many proofs of inference, becoming inefficient in a ZKP system.
@@
Our algorithm samples $m$ directions and then starting from the original input $x$, takes $\delta$ steps until it finds a point with a different label along all these directions individually. The border point $x_{border}$ is that oppositely labeled point which is closest to the input $x$. Furthermore, unlike in the algorithm in (Laugel et al., 2018a), our algorithm can exploit parallelization by searching along the different directions in parallel since these are independent.
@@ 8 Related Work
In the ML field ZKPs have been majorly used for verification of inferences made by models (Sun et al., 2024; Chen et al., 2024; Kang et al., 2022; Weng et al., 2023; Sun & Zhang, 2023; Feng et al., 2021; VI2, 2023; Lee et al., 2020; Zhang et al., 2020a; Liu et al., 2021; Singh et al., 2022; Fan et al., 2023). A line of work also focuses on proving the training of ML models using ZKPs (Lycklama et al., 2023; Huang et al., 2022; Rückel et al., 2022; Garg et al., 2023; Abbaszadeh et al., 2024). More recently they’re also been used for verifying properties such as fairness (Yadav et al., 2024a; Shamsabadi et al., 2023; Toreini et al., 2023) and accuracy (Zhang et al., 2020b) of confidential ML models. Alternatively, (Waiwitlikhit et al., 2024) propose ZKPs as a general technique for auditing ML models broadly. Contrary to these and to the best of our knowledge, ours is the first work that identifies the need for proving explanations and provides ZKP based solutions for the same.